1.) 4u=37+4u
2.) 5x+8=7x+8
3.) 2x-5=3x+4

Answer by user_dude2008(1862)

1.) 4u=37+4u
4u-4u=37
0=37
no solution
2.) 5x+8=7x+8

5x-7x=8-8
-2x=0
x=0
3.) 2x-5=3x+4

2x-3x=4+5
-x=9
x=-9